OpenStudy (anonymous):

hair grows at a rate of 1/64 inch per day write this number using negative exponents

OpenStudy (mathstudent55):

\(64 = 2^6\), so \(\dfrac{1}{64} = \dfrac{1}{2^6} = 2^{-6} \)